OpenStudy (dilla):

when solving an equation, how do you determine what number to add,subtract,multiply or divide by on both sides of that equation?

OpenStudy (dtan5457):

It you mean when your trying to solve for x or y. The point is to isolate the variable.

OpenStudy (dtan5457):

Hmm example..

OpenStudy (dtan5457):

3x+6=3 ISOLATE variable. Subtract the 6. 3x=3-6 3x=-3 3 is front of the x, and it's multiplication. Therefore you divide by 3. x=-1

OpenStudy (dilla):

That's all the question asks. I just don't know how to word it. Best I can get is.......You have to get the variable alone so you do the opposite of the equation. ex: If it's x+3=6 ,you do the opposite of +3 which is -3.

OpenStudy (dtan5457):

Sounds about right

OpenStudy (dtan5457):

"isolate the variable" is the important part

OpenStudy (dilla):

Thank you so much!